# Consistency of Responses and Continuations Generated by Large Language Models on Social Media

### v2

@@ Results
Reddit anger 197.316 6.55E-166
anticipation 80.899 1.18E-67
disgust 35.307 4.12E-29
optimism 59.526 6.83E-50
fear 45.168 2.70E-37
sadness 18.736 3.58E-15
Twitter anger 384.285 9.5078E-320
anticipation 16092.479 0
disgust 179.181 1.80E-143
optimism 6.088 6.95E-05
fear 73.936 6.99E-60
joy 1742.163 0.00E+00
sadness 148.411 6.85E-98
Analysis of variance (ANOVA) results presented in Table 11 indicate statistically significant differences ($P<0.05$) across the five groups in emotional intensity values for anger, anticipation, disgust, optimism, fear, joy, and sadness on Twitter. Similar significant variations were observed in the Reddit dataset for anger, anticipation, disgust, optimism, fear, and sadness. Tukey’s post-hoc analysis identified several significant differences across three comparison categories: within-model, between-model, and model-to-human comparisons.
On Twitter discussions of climate change, within-model analyses revealed that Gemma’s continuation content exhibited significantly lower anticipation values compared to its response content. Similarly, Llama’s continuation content showed significantly lower anticipation values than its response content.
Between-model comparisons demonstrated that Gemma’s continuation content expressed significantly higher anticipation values than Llama’s continuation content. Additionally, Gemma’s response content showed significantly higher anticipation values compared to Llama’s responses. In model-to-human comparisons, both models’ generated content (continuation and response) demonstrated significantly lower anticipation values compared to the original human text.
These findings demonstrate that: (1) both models express higher anticipation values in response tasks compared to continuation tasks; (2) Gemma consistently generates content with higher anticipation values compared to Llama across both tasks; and (3) both models generate content with significantly reduced anticipation values compared to human-authored text, indicating a systematic reduction in emotional intensity during the generation process.
Table 2: Tukey’s post-hoc test of Reddit emotions
Emotion Within Group1 Between Groups2 Comparison with original
Gemma Con
vs. Resp
Llama Con
vs. Resp
Gemma Con
vs. Llama Con
Gemma Resp
vs. Llama Resp
Gemma Con
vs Original
Gemma Resp
vs Original
Llama Con
vs. Original
Llama Resp
vs. Original
Anticipate <*** <*** <*** <***
Joy <**
Disgust <*** <** >*** <*** >* <***
Sadness <*** >* <*** <*** <***
Anger <*** <*** <*** <*** <*** >***
Optimism >*** >** <*** <*** <*** <***
Fear <*** <*** <*** <*** <***
@@
    The symbolic $<$ and $>$ indicate that the data in the previous column is less than or greater than the data in the next column. The “-” means there is no statistically significant difference between the two columns of data.
Table 3: Tukey’s post-hoc test of Twitter emotions
Emotion Within Group1 Between Groups1 Comparison with original1
Gemma Con
vs. Resp
Llama Con
vs. Resp
Gemma Con
vs. Llama Con
Gemma Resp
vs. Llama Resp
Gemma Con
vs Original
Gemma Resp
vs Original
Llama Con
vs. Original
Llama Resp
vs. Original
Anticipate <*** <*** >*** <*** <*** <*** <*** <***
Joy <*** <*** <*** <***
Disgust <*** <*** <* <*** <***
Sadness <*** <*** <*** <*** <*** <***
Anger <*** <*** <*** <*** <*** <*** <*** <***
Optimism <*** <*** <*** <***
Fear <*** <*** <** <** <*** <*** <*** <***
@@
  • 3
@@
    The symbolic “$<$ and $>$” indicate that the data in the previous list is less than or greater than the data in the next list. The “-” means there is no statistically significant difference between the two columns of data.
Further analysis of additional emotional dimensions revealed that Tukey’s post hoc test results indicate significant differences ($P<0.05$) between LLM-generated and original texts across multiple emotional categories. Both Gemma and Llama’s generated content, in both continuation and response tasks, exhibited significantly lower intensities of positive emotions, particularly joy and optimism. A nuanced pattern emerged in the expression of negative emotions. The continuation texts generated by both models demonstrated attenuated levels of sadness, anger, disgust, and fear compared to their respective response texts and the original human content. Notably, Gemma’s response text exhibited significantly lower intensities of anger, disgust, and fear compared to Llama’s responses. These findings suggest two key insights: first, LLMs demonstrate systematic suppression of certain negative emotions, particularly in continuation tasks; second, the response task appears to operate under distinct generative mechanisms, resulting in differential emotional expression patterns. Furthermore, the consistent reduction in optimism across all LLM-generated texts relative to human-authored content indicates a systematic constraint in LLMs’ capability to fully capture and convey positive emotional states.
Semantic Coherence Analysis between LLM-Generated and Human-Authored Content
In LLM-human interactions, beyond examining emotional congruence and intensity patterns demonstrated in our previous experiments, we investigated the models’ capacity to maintain topical coherence and generate contextually relevant responses. To quantify this relationship, we employed cosine similarity as a metric to assess semantic alignment between generated and original content. This approach provides a quantitative framework for evaluating the semantic fidelity of LLM-generated responses across different interaction contexts.
@@
Figure 6: Distribution of cosine similarity scores for LLM-generated content, visualized using violin plots comparing Gemma and Llama models’ continuations and responses across Reddit (panel a) and Twitter (panel b) platforms. Cosine similarity values range from -0.2 to 1.0, with median scores indicated numerically for each distribution.
Table 4: Mann-Whitney U test for cosine similarity between LLM-generated content and two platform text
Comparison
group
Statistics P value
Reddit Inner
Gemma
continuation
vs
Gemma
response
56740052 0.999999995
Llama
continuation
vs
Llama
response
66851679 4.05E-58
Inter
Gemma
continuation vs
Llama
continuation
68723025 8.90041E-90
Gemma
response vs
Llama
response
78292192 0.00E+00
Twitter Inner
Gemma
continuation vs
Gemma
response
93014920 9.74E-251
Llama
continuation vs
Llama
response
60081284 1
Inter
Gemma
continuation vs
Llama
continuation
118641175 0
Gemma
response vs
Llama
response
69518835 1
  • 1
    The $p<0.05$ indicate that the data in the previous group is significantly higher than the data in the latter group.
Table 4 presents the Mann-Whitney U test results across the four experimental conditions. The analysis reveals that both Gemma and Llama maintained substantial semantic alignment with the original text in their continuation and response tasks. Figure 6a illustrates that semantic similarity values on the Twitter platform predominantly exceeded 0.5, indicating strong semantic coherence between LLM-generated content and input text. This suggests the models’ capability to comprehend and generate contextually relevant responses while maintaining topical coherence. Similar patterns emerged in the Reddit dataset, as shown in Figure 6b, where all four sets of cosine similarities demonstrated values above 0.5.
The Mann-Whitney U test (Table 4) revealed significant within-model and cross-task variations in semantic similarity. On the Twitter platform, Gemma’s continuation task demonstrated significantly higher semantic similarity compared to its response task, while Llama exhibited the opposite pattern, with responses showing significantly higher similarity ($P<0.01$). Cross-model analysis revealed task-specific differences: Gemma’s continuation task demonstrated significantly higher semantic similarity compared to Llama’s continuation task ($P<0.01$), while Llama achieved higher similarity scores in the response task. On Reddit, Llama’s continuation content exhibited significantly higher semantic similarity compared to its response content ($P<0.01$). Furthermore, Gemma’s generated content demonstrated significantly higher semantic similarity compared to Llama’s in both continuation and response tasks ($P<0.01$).
@@ Discussion
This study investigated the consistency between LLM-generated content and human input, with particular focus on emotional content transmission. Given that text often encodes substantial emotional information, we hypothesized that LLMs may demonstrate diverse emotional patterns. As LLMs increasingly integrate into human social contexts through chatbots and social agents, understanding their semantic consistency and emotional transmission patterns becomes crucial for human-machine interaction. We decomposed the fundamental question of LLM emotional capability into three specific research questions, based on our empirical findings. Several key insights emerged from our analysis.
In the continuation task, Gemma demonstrated a systematic transformation of most original emotions toward anger, indicating a bias toward negative emotional expression, particularly in anger intensification. However, Gemma exhibited capacity for emotional preservation, specifically for optimism and surprise, suggesting ability to perceive and maintain certain emotional valences through the continuation process. Conversely, the Llama model demonstrated enhanced capability in preserving original emotional states during continuation tasks, particularly for anger, anticipation, fear, optimism, and sadness, with relatively lower emotional transformation rates, indicating superior emotional continuity (RQ1).
We posit that during continuation tasks, LLMs employ emotion-specific lexical markers, syntactic patterns, and contextual cues to maintain emotional consistency with source texts. Emotions characterized by prominent semantic features (such as anger, anticipation, and sadness) typically manifest through distinct emotional vocabulary and strong affective markers, facilitating model recognition and continuation. Moreover, the semantic salience of anger may enhance its recognition and preservation in generative tasks, while positive emotions like optimism or joy may require more robust contextual support, making them susceptible to disruption by LLM-generated negative content.
In the response task, both Gemma and Llama models exhibited a systematic bias toward positive emotions, predominantly converting original emotions into anticipation and optimism. This pattern suggests an embedded bias from the training process favoring positive affect. However, the persistent presence of anger in a subset of responses demonstrates these models’ susceptibility to negative emotional content during the response generation process.
Our experimental findings indicate that both Gemma and Llama models exhibit emotional capabilities and can effectively recognize and maintain human emotional states in continuation tasks. This emotional recognition capability is fundamental for LLMs, as it underlies their ability to comprehend both user requirements and emotional context.
Regarding emotional intensity (RQ2), our results demonstrate that LLM-generated content exhibits significantly attenuated intensity across most emotional dimensions compared to human-authored texts. This suggests that LLMs exhibit more moderated emotional expression, maintaining more constrained emotional ranges across both positive and negative affects. Furthermore, the consistently lower emotional intensity values in both continuation and response tasks, compared to original texts, indicates potential limitations in LLMs’ capacity to fully capture and convey emotional depth.
Analysis of emotional types reveals distinct patterns in LLM-generated content. Regarding positive emotions, LLM-generated texts demonstrate significantly lower intensities of joy and optimism compared to human-authored texts, suggesting limitations in the models’ capacity to fully comprehend and express positive emotional states characteristic of human experience. Similarly, for negative emotions, LLM-generated content exhibits attenuated intensities of sadness, anger, disgust, and fear, indicating systematic suppression of negative emotional expression in downstream tasks.
@@
In the context of global challenges such as climate change, which transcends temporal, spatial, and geographical boundaries and impacts long-term human development, public discourse exhibits diverse emotional responses. As social media platforms rapidly evolve, they become crucial channels for public communication during climate-related events, such as the Australian bushfires. The effective management of public emotions during critical periods following such events represents a key component of public opinion governance. Recent advances in Artificial Intelligence Generated Content (AIGC), driven by Generative AI (GAI) technology, have garnered attention beyond computer science (Cao et al. 2023). Given the increasing integration of LLMs into daily life, their emotional characteristics significantly influence opinion leadership, as emotional content shapes public perception and discourse framing.
Our experimental results (Figure 3 and Figure 4) demonstrate that LLMs systematically transition toward positive emotions during response tasks. This positive bias suggests potential utility in public opinion events, offering constructive responses and mitigating negative public sentiment. However, the observed patterns of negative emotional transmission pose potential societal risks. LLMs’ capability to generate contextually relevant content could be exploited for malicious purposes. Beyond concerns regarding misinformation dissemination, the potential intensification of emotional polarization presents a significant challenge in LLM deployment.
From the perspective of human-computer interaction, emotional support serves as a fundamental component in enhancing social interactions, facilitating psychological interventions, and improving customer service outcomes through addressing emotional needs. The quality of emotional support and user understanding significantly impacts long-term user engagement and trust in LLM interactions (Schneider, Flores, and Kranz 2024).
Addressing our third research question, we analyzed embedding representations of generated and original content, employing cosine similarity as a metric for semantic proximity. Our findings indicate that both Gemma and Llama consistently demonstrate high cosine similarity values, suggesting robust capability in capturing and reproducing semantic features. Models have shown a remarkable ability to represent, comprehend, and generate human-like text. Compared to prior Natural Language Processing (NLP) approaches, one of the most striking advances of LLMs is their ability to generalize their “knowledge” to novel scenarios, contexts, and tasks(Peters and Matz 2024) Future implications suggest that while LLMs demonstrate competence in generating semantically coherent content, opportunities exist for enhancing alignment between generated outputs and nuanced human context. Future research directions should explore mechanisms for refining LLMs’ comprehension of implicit meaning and contextual subtleties, thereby enhancing user experience and expanding application domains.
@@
This study contributes to understanding the emotional dynamics of human-AI interactions, while suggesting several avenues for future research based on current limitations.
First, regarding experimental design, our study was limited to data from Reddit and Twitter platforms, potentially under-representing the broader social media ecosystem. Other platforms such as YouTube, Instagram, and TikTok exhibit distinct user behaviors and content structures (Hilde A. M. Voorveld and Bronner 2018). Furthermore, our analysis was bounded by the capabilities of open-source emotion models (Llama and Gemma), which may demonstrate different performance characteristics compared to proprietary commercial models like ChatGPT or Google’s Bard. Future research could expand the scope by incorporating data from additional social media platforms and evaluating commercial models to enhance generalizability.
Second, our experimental scope, focused on English-language climate change discourse, provides limited insight into model performance across diverse topics and languages. The emotional expression patterns of LLMs may vary significantly across different subject domains. Future research directions could incorporate multilingual datasets to identify cross-linguistic semantic variations (Zhao et al. 2020) and investigate potential biases within LLMs across isolated or intersecting semantic spaces (Hassan et al. 2018).
Third, while we utilized social media to obtain human-authored content, the dataset potentially includes automatically generated content and social bot interactions that may not accurately represent genuine human expression. Future research should implement more rigorous data validation protocols to enhance dataset quality and control for confounding variables.

### v3

@@ Results
Reddit anger 114.2381 <0.001
joy 80.899 <0.001
disgust 29.8564 <0.001
surprise 24.9967 <0.001
fear 17.3482 <0.001
sadness 5.5149 <0.001
neutral 325.2502 <0.001
Twitter anger 202.1090 <0.001
joy 34.4011 <0.001
disgust 22.4241 <0.001
surprise 51.5744 <0.001
fear 75.0905 <0.001
sadness 9.2834 <0.001
neutral 552.6669 <0.001
Analysis of variance (ANOVA) results presented in Table 1 indicate statistically significant differences ($P<0.01$) across the nine groups in emotional intensity values for anger, disgust, fear, sadness, joy, surprise and neutral on Twitter. Similar significant variations were also observed in the Reddit dataset for the seven emotions mentioned above. Tukey’s post-hoc analysis identified several significant differences across three comparison categories: within-model, between-model, and model-to-human comparisons.
Each emotion score represents its degree of emotion, and the higher the score, the stronger the emotion. As shown in the table2, intra-model comparisons indicate that emotional scores for certain emotions—such as anger and fear—are generally higher in continuation tasks than in response tasks. This suggests that LLMs tend to adopt a more neutral emotional tone when operating as conversational agents. Furthermore, when compared to the original human-authored texts,shown in the table3, most LLM-generated outputs exhibit lower emotional intensity, particularly in dimensions such as anger, fear, disgust, and joy. This may indicate that LLMs express emotions in a more restrained and subdued manner relative to humans.
Cross-model comparisons also reveal that emotional expression varies significantly depending on task type and data source. For instance, within the Llama model family, Llama 3 demonstrates higher emotional scores than Llama 3.3, suggesting a greater tendency toward heightened emotional expression.
Table 2: Tukey’s post-hoc test of LLM-generated content emotion values
Emotion Within Group1 Between Groups2 Platform
Gemma Con vs Resp Llama Con vs Resp Claude Con vs Resp Llama3.3 Con vs Resp Gemma Con vs Llama Con Gemma Con vs Claude Con Llama Con vs Claude Con Llama Con vs Llama3.3 Con Gemma Resp vs Llama Resp Gemma Resp vs Claude Resp Llama Resp vs Claude Resp Llama Resp vs Llama3.3 Resp
anger >*** <*** >*** - >** - >*** - <*** <* <*** >*** Reddit
fear <*** - <** <*** <*** >*** - >* - - - -
sadness >** <*** <** - >*** - >*** - <** >*** >*** -
joy - - >*** >*** >*** - >*** - >* <** - >***
surprise - >** - - >*** <** - - >*** <* - -
disgust <* <*** <** <*** - - - >*** - - - -
neutral >*** >*** >** >*** - <*** <*** >*** <*** >* <*** >***
anger - - >*** >*** >*** - <*** - >* >** - >*** Twitter
fear >*** >*** >** >*** - >*** >*** >*** <*** <*** >*** >***
sadness <*** - <* <*** <*** <*** - >* - - - -
joy - >*** - - >*** >** - - >*** >* - -
surprise >* <*** <** - >*** - <*** - <** <*** <*** -
disgust >*** <*** >*** - >** - <*** - <*** >* >*** >***
neutral <* <*** <** <*** - - - >*** - - - -
@@
    The symbolic $<$ and $>$ indicate that the data in the previous column is less than or greater than the data in the next column. The “-” means there is no statistically
    significant difference between the two columns of data.
Table 3: Tukey’s post-hoc test of LLM-generated content and human text emotion values
Platform Emotions Gemma Con vs Original1 Gemma Resp vs Original Llama Con vs. Original Llama Resp vs. Original Claude Con vs Original Claude Resp vs Original Llama3.3 Con vs Original Llama3.3 Resp vs Original
Reddit anger <*** <*** <*** <*** <*** <*** <*** <***
fear <*** - - - - - <*** -
sadness - - - - >* >*** - -
joy - - <*** - - <** <** <***
surprise - - <*** <*** - - <*** -
disgust >*** >*** >*** >*** >** >*** - >***
neutral >*** <*** >*** <** <*** <*** - <***
Twitter anger - - <*** - - <** <** <***
fear >*** <*** >*** <** <*** <*** - <***
sadness <*** - - - - - <*** -
joy - - <*** <*** - - <*** -
surprise - - - - >* >*** - -
disgust <*** <*** <*** <*** <*** <*** <*** <***
neutral >*** >*** >*** >*** >** >*** - >***
@@
    The comparison of the emotion values of the LLM-generated content by different model and human text.
@@
    The stars indicate the p values of the Tukey’s post-hoc test: *** for $p<0.001$, ** for $p<0.01$, and * for $p<0.05$.
@@
    The symbolic “$<$ and $>$” indicate that the data in the previous list is less than or greater than the data in the next list. The “-” means there is no statistically
    significant difference between the two columns of data.
These findings suggest two key insights: first, LLMs demonstrate systematic suppression of certain negative emotions, particularly in continuation tasks; second, the response task appears to operate under distinct generative mechanisms, resulting in differential emotional expression patterns. Furthermore, the consistent reduction in optimism across all LLM-generated texts relative to human-authored content indicates a systematic constraint in LLMs’ capability to fully capture and convey positive emotional states.
Evaluating Semantic Consistency of LLM-Generated Content in Social Media Contexts
In LLM-human interactions, we analyzed models’ ability to maintain topical coherence and contextual relevance using scores labeled by LLM-as-judge (Zheng et al. 2023) to measure semantic and logic alignment with the original content, providing a framework for evaluating semantic and logic fidelity across contexts.
@@
Figure 6: Distribution of faithfulness and consistency scores for LLM-generated content, visualized using stacked histograms comparing Gemma, Llama, and Claude models’ continuations and responses across Reddit (panel a) and Twitter (panel b) platforms. The x-axis denotes rating categories, with 5 representing ’Excellent’, 4 ’Good’, 3 ’Fair’, 2 ’Poor’, and 1 ’Very Poor’.
Figure 6 plotted the frequency of each score (ranging from 1 to 5) separately for each task. This allows us to examine whether certain tasks tend to produce more consistent model outputs. The results suggest that while most tasks concentrate around higher scores, there are subtle differences in how each subtask is rated by the evaluator. We find that most models are capable of generating high-quality content in both response and continuation tasks, indicating that current LLMs possess strong capabilities in understanding and interacte with human language. However, subtle differences remain across models. For instance, the Llama 3.3 model demonstrates consistently strong performance in content generation across both types of datasets. Comparized with Llama 3’ result, this may also indicate to some extent the impact of different model sizes on the quality of generated content. Additionally, model performance varies across platforms. For example, the Claude model shows higher fidelity and consistency in its outputs on the Reddit dataset, which may be attributed to the more complex linguistic environment of Twitter.
@@ Discussion
This study utilized the Twitter and Reddit datasets on climate change to systematically evaluate three aspects of LLM-generated text versus original social media posts: (1) emotional consistency in both continuation and response tasks (and differences in different models), (2) emotional intensity across those tasks, and (3) logic similarity and context-awareness between model outputs and original social media texts. By combining these analyses, we aim to uncover the potential effects on emotional dynamics that LLMs have when implemented in online discussion.
In the continuation task, Gemma consistently transformed most emotions toward anger, showing a bias toward negative expression and intensifying anger. However, it maintained optimism and surprise, suggesting some ability to preserve emotional valence. On the other hand, Llama better preserved original emotions like anger, anticipation, fear, optimism, and sadness, with fewer emotional shifts, reflecting stronger emotional continuity (RQ1).
By combining the comparative analyses of emotional consistency and intensity, this study demonstrates that in both response and continuation tasks, LLMs systematically shift emotional content toward neutrality and lower its intensity across most emotional categories compared to original texts. Particularly, generated outputs display a higher prevalence and intensity of “neutral” manner, especially in response tasks rather than continuation tasks. This pattern shows that LLMs do have effect on the change of emotion when engaged in online discussion, and they’re turning the discussion in a emotionally-moderate way.
As aforementioned, emotion plays a critical role in navigating social media discussion, especially in controversial topics like climate change. While the climate change discussion online is quite polarized (Treen et al. 2022b), such a tendency to neutrality could help defuse heated exchanges and foster calmer, less contentious dialogue. Hence, even though LLMs cannot keep the emotional consistency as human, its preferences toward moderation could benefit the highly polarized and emotionally debatable topics online, and lead the discussion more neutrally.
However, beyond this “neutral” tendency, what should be aware is that the negative emotions like “anger” and “fear” also showed increase in LLM-generated continuations and responses in most of the cases, which indicates that these models can amplify such emotional cues. Anger and fear are among the most salient emotional frames in climate discourse (Nabi, Gustafson, and Jensen 2018; Davidson and Kecinski 2022), and are significantly associated with climate activism, engagement, and has great effect on people’s attitudes and behaviors (Stanley et al. 2021; Gregersen, Andersen, and Tvinnereim 2023). An uptick in fear and anger in LLM outputs may therefore shape users’ perceptions of climate change. Considering the echo chamber effect on social media, LLMs’ propensity of fear and anger could increase people’s perceived threat by climate change, which in turn leads to higher awareness and action intentions towards climate crisis.
In terms of semantic consistency, the results of LLM-as-judge (Figure 6) show that LLMs generated text could generally keep the consistency to social media texts. This comparative findings proves a critical evidence from the perspective of emotion that LLMs could generally “understand” the context of a social media post and produce continuations or replies that align both thematically and logically with the source content. However, the consistency varies due to different types of LLMs, while most of the scores are high, some specific model like Claude’s performances are dispersed, featuring a nontrivial tail into other scores. It indicates that some models may be more sensitive to prompt phrasing or exhibit less robust world-knowledge integration. The findings from semantic consistency imply that LLMs nowadays can generally preserve semantic and logic coherence in social-media dialogue, which makes the LLMs-generated content hard to distinguish on social media.
Taken together, the emotional and semantic consistency examined in this study could be seen as a technological improvement of context-awareness in LLMs’ text generation capabilities. However, it could also raise practical concerns to general social media users for the risks of LLMs being manipulated for some malicious or intentional purposes, such as misinformation.
@@
Effectively managing public emotions during controversial discussion online is crucial for governance of public opinion. Recent advances in Artificial Intelligence Generated Content (AIGC), driven by Generative AI (GAI) technology, have garnered attention beyond computer science (Cao et al. 2023). Given the increasing integration of LLMs into daily life, their emotional characteristics significantly influence opinion leadership, as emotional content shapes public perception and discourse framing.
Future implications suggest that while LLMs generate semantically coherent content, there is potential to improve alignment with nuanced human contexts. Research should focus on refining LLMs’ understanding of implicit meaning and contextual subtleties to enhance user experience and broaden application domains.
@@
This study enhances understanding of emotional dynamics in human-AI interactions while highlighting opportunities for future research to address current limitations.
First, regarding experimental design, our analysis relied on Reddit and Twitter data. Platforms like YouTube, Instagram, and TikTok exhibit unique user behaviors and content structures (Hilde A. M. Voorveld and Bronner 2018).
Second, the explaination for the findings in emotional inconsistency of LLMs requires further exploration. While we revealed that emotions vary both across different LLMs and across generation tasks, for example, the “neutral” emotion take the major role among all emotions in all cases but differs between continuation and response tasks and LLMs, likewise, the overall proportion of “anger” significantly increases in all cases, and different models exhibit distinct emotional profiles. However, because the training data and internal weights of these models are not publicly available, we cannot at present pinpoint the underlying causes of these differences—whether they stem from pretraining data biases, fine-tuning strategies, or discussion context. Future work should work on the interpretability analyses, controlled comparisons, or replication on open datasets, to investigate the mechanisms driving these emotion changes.
Third, this study may have potential data contamination due to the use of publicly available data from Twitter and Reddit. Twitter and Reddit data are publicly accessible and may overlap with the LLMs’ pre-training corpora, model outputs could partly reflect memorized patterns rather than genuine reasoning in real-world contexts, which limits the generalization of the results. While we cannot fully quantify the contamination, we acknowledge its impact and encourage future research to explore methods for mitigation—such as by using post-training data that is demonstrably out-of-scope of LLM training sets, or by focusing on data published after known pretraining cutoff dates. Besides, social media-derived content may include automated posts and social bot interactions, potentially skewing results. Future research should adopt stricter data validation protocols to ensure datasets quality and reduce confounding variables.
Fourth, since human expert evaluation is expensive, we adopt a hybrid evaluation strategy that combines human judgment with LLM assistance based on a small but diverse sample across platforms, tasks, and models. While this approach achieves high-quality and interpretable scores, we acknowledge that the limited sample size may not capture the full distribution of consensus behaviors across the datasets. Future work could extend this approach, such as scalable pure LLM evaluation and automatic divergence detection for selective human review.